OpenStudy (anonymous):

A line passes through the point (0, 2) and has a slope of -1/4. What is the equation of the line? A. 4x+y=2 B. x-4y=8 C. 4x-y=-2 D. x+4y=8

OpenStudy (texaschic101):

y = mx + b slope(m) = -1/4 (0,2)...x = 0 and y = 2 now we sub.....and solve for b, the y intercept 2 = -1/4(0) + b 2 = b equation is : y = -1/4x + 2...but we need it in standard form Ax + By = C y = -1/4x + 2 ---- multiply entire equation by 4 4y = -x + 8 --- add x to both sides x + 4y = 8

OpenStudy (anonymous):

thank you! i finally understand how to do problems like that! the way you explained it actually makes since. so the answer would be "D"